22. What is the pH of 0.12 M HCIO (aq) (Ka = 3.0 x 10)?
Answer #1

22) Ka= 3x10-8 concentration - Co= 0.12 M considering Reaction - - Helo199) Ť Heag + c10 199) Initialy. Co o o At : coll-a) cnow, Check the value of ko Case Ii 97 kg is greater than 0.1 then solve the quadratic equation (equestion) to find the value[4+] = 6xlots om pH = -log {sot} = -log (6x10-5 M) 25 - logo -4.22 pH = 4.22 T AwypH=4.22